Two mettalic wires A and B are connected  in second wire A has length l and radius r , while wire B has length 2l amd radius 2r. Find the ratio of total resistence of series combination and the resistence of wire A, if both the wires are of same material ?
 

Grade:10

1 Answers

Vikas TU
6 years ago
Suppose the resistivity of both the wires is ‘q’.
Resistance of wire A is, R = ql/(πr )
Resistance of wire B is, R = q(2l)/[π(2r)
Now, the resistance when they are connected in So, the required ratio is, R: R = (3/2)[ql/(πr)]
=> R: R = 3: 2